OpenStudy (anonymous):

What is the distance between (2, 0) and (5, 8)? I'm horrible at the whole distance thing!

jimthompson5910 (jim_thompson5910):

d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((5-2)^2+( 8- 0)^2) d = sqrt((3)^2+(8)^2) d = sqrt(9+64) d = sqrt(73) So exact distance between the two points is sqrt(73) units.

OpenStudy (anonymous):

thanks soo much jim_thompson5910!